OpenStudy (anonymous):

What would this absolute value equation look like on a graph ? HELP , pleaseee ? y = –1/2 |x|

OpenStudy (anonymous):



OpenStudy (anonymous):

thank youuu soo much (;

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

can you help me with this next one ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

y = 7|x + 8| –4 y = –7|x – 8| +4 y = –7|x + 8| –4 y = –7|x – 8| –4

OpenStudy (anonymous):

which function matches that graph? ^

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thank youuu (:

OpenStudy (anonymous):

yw